Can you show me step by step how to solve this equation?
You deposit $1000 into a savings account in which the intrest is compounded anuualy at a rate of 5%.

What is the pricipal after 5 years?

Found 2 solutions by mananth, ikleyn:
Answer by mananth(16949)

Find Amount
Principal P = 1000
Amount= A
years=n 5.00
compounded 1,times a year t
Rate = 5.00,=0.05
Amount = P*((n+r)/n)^n*t
Amount = 1000*(1+ 0.05)^5* 1
Amount = 1000*(1+ 0.05)^5
1000*(1.05)^5*1
Amount = $1276.28

Answer by ikleyn(53750)

.
Can you show me step by step how to solve this equation?
You deposit $1000 into a savings account in which the intrest is compounded anuualy at a rate of 5%.

What is the pricipal after 5 years?
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~


The question in the post is posed incorrectly and shows
that a person who created/composed the problem did not know the terminology of the subject.

A correct question should ask about the Future value of the account in 5 years.

The principal is and remains unchangeable.